1. Supposed I were to tell you that there are 5.2 million white females with diabetes in the United States 20 years and older out of a population of 84.7 million, and 2.2 million black females with diabetes out of the population of 14.0 million. what is the ratio (white to black) of females with diabetes? what does this tells you? what does it not tell you?

2. Out of all white women age 20 and above, what proportion of them have diabetes? what does this tells you? what does it not tell you?
Answer #1

1.In united states of 20 years and above having diabetes,comparing between white females and black females,

Ratio=5.2 million white females of population 84.7 million/2.2 million black females of population 14.0 million

Ratio =5.2×14/84.7×2.2

Ratio=75.6/186.34 or 75.6:186.34

*It tells that, we can calculate the ratio of white and black females with diabetes of 20 years and above in the united states.*it can calculate RISK RATIO of white and black females with diabetes.

*But,we can't calculate the individual ratio of white females or individual ratio of black females with diabetes of 20 years and above in the united states.

2.In white women of 20 years and above in the united states,

Proportion=Number of white females with diabetes/total number of population in white females×100

Proportion=5.2/84.7×100

Proportion=0.0613×100

Proportion of white females=6.13%.

*It tell that Proportion of white females with diabetes individually can be calculated,finally tell the INCIDENCE RATE of diabetes in white females.

*But,it doesn't calculate the proportion of both white and black females with diabetes in one equation.